# FRoundation: Are Foundation Models Ready for Face Recognition?

### v2

@@ Abstract
Foundation models are predominantly trained in an unsupervised or self-supervised manner on highly diverse and large-scale datasets, making them broadly applicable to various downstream tasks. In this work, we investigate for the first time whether such models are suitable for the specific domain of face recognition. We further propose and demonstrate the adaptation of these models for face recognition across different levels of data availability. Extensive experiments are conducted on multiple foundation models and datasets of varying scales for training and fine-tuning, with evaluation on a wide range of benchmarks. Our results indicate that, despite their versatility, pre-trained foundation models underperform in face recognition compared to similar architectures trained specifically for this task. However, fine-tuning foundation models yields promising results, often surpassing models trained from scratch when training data is limited. Even with access to large-scale face recognition training datasets, fine-tuned foundation models perform comparably to models trained from scratch, but with lower training computational costs and without relying on the assumption of extensive data availability. Our analysis also explores bias in face recognition, with slightly higher bias observed in some settings when using foundation models.
@@ 1 Introduction
Significant progress has been made in developing foundation models trained on extensive and diverse datasets recently [33, 2, 23]. Once these models are trained, they can be adapted to a wide array of downstream tasks, providing a versatile basis. Inspired by the success of large language models (LLM) [18, 8, 39, 13, 38], similar large-scale foundation models have been explored for various perception tasks [35, 33, 2, 23]. Often based on the ViT architecture [19], which has been shown to match or exceed the performance of traditional methods in large-scale image classification tasks, visual foundation models are becoming increasingly popular due to their strong generalization capabilities when trained on large datasets. Although foundation models hold considerable promise for a wide range of applications, their adaptation for face recognition has not been explored in previous research, which drives the motivation for this study.
@@
To explore the potential of foundation models in face recognition, we evaluate the performance of different versions of two widely used foundation models, namely DINOv2 [33] and CLIP [35]. Although foundation models demonstrate strong generalization capabilities, our experiments show that they perform poorly compared to state-of-the-art (SOTA) face recognition models [17, 42, 4, 16] on various benchmarks. To enhance their effectiveness for the downstream task of face recognition, we propose to fine-tune the considered foundation models using low-rank adaptation (LoRA) [24]. LoRA integrates trainable low-rank decomposition matrices into each transformer block while keeping the pre-trained model weights frozen. After fine-tuning on dedicated datasets, the LoRA layers adapt to the downstream task of face recognition, resulting in a significant increase in performance. For example, the smallest released pretrained models of DINOv2 and CLIP achieve average verification accuracies of 64.70% and 82.64%, respectively. After fine-tuning on Casia-Webface, their accuracies increase to 90.94% and 92.13%, respectively.
In addition to examining the performance of fine-tuned foundation models, we compare them to a Vision Transformer (ViT) trained from scratch on different subset sizes of data. This comparison validates the advantages of using foundation models over training an entire model based on data availability. Finally, we investigate the bias in face recognition and demonstrate that it tends to be slightly higher when fine-tuning foundation models compared to training from scratch.
@@ 2 Related Work
Foundation models are defined by a substantial number of trainable parameters and are pre-trained on a large and diverse dataset, enabling them to adapt to a wide range of downstream tasks across various domains [3]. Vision foundation models are commonly structured around the use of Vision Transformers (ViTs) [19] and rely on self-supervised learning (SSL) [21]. SSL is a technique that trains models to learn representations from unlabeled data and is essential for training ViTs, which tend to perform poorly on small datasets [54]. Several high-performing vision foundation models specialized in various tasks have been released in recent years. Building on the success of the Segment Anything Model (SAM) [28], which introduced a foundation model for object segmentation, SAM2 [36] presents a model for real-time, promptable object segmentation in images and videos, achieving state-of-the-art performance. CLIP [35] learns visual concepts from natural language supervision and can be applied to any visual classification by providing the names of the visual categories to be recognized. DINOv2 [33] foundation models generate universal features suitable for both image-level visual tasks, such as image classification and video understanding, as well as pixel-level visual tasks, including depth estimation and semantic segmentation.
@@
Vision foundation models are characterized by their generalization ability due to massive training data, but they tend to show poor performance when applied to domain-specific settings [41]. To adapt them to a downstream task, multiple approaches were considered in the literature for this adaption. An example of that is the AdaptFormer [11] that replaces the MLP block in the transformer encoder with two identical MLP branches, where one mirrors the original network and the other introduces a lightweight module for task-specific fine-tuning, demonstrating significant improvements compared to fully fine-tuned models on downstream tasks. Another example is the ViT-Adapter [12] that proposes a technique to adapt plain ViTs for dense prediction tasks by injecting image-related inductive biases into the ViT and reconstructing fine-grained multi-scale features, yielding SOTA results on COCO test-dev. Another approach is to insert trainable rank decomposition matrices, called Low-rank adaptation (LoRA) [24] layers, while freezing the pre-trained model weights. In this work, we choose LoRA as our foundation model adapter, as recent studies [49, 15, 10, 48, 41] highlight the significant potential of LoRA for this purpose. For instance, integrating LoRA layers into DINOv2 has been successfully applied in the medical domain for two distinct tasks: capsule endoscopy diagnosis [49] and depth estimation in endoscopic surgery [15], both of which have demonstrated superior performance in their respective fields. In another work on adapting foundation models for multiple plant phenotyping [10], Chen et al. demonstrated that LoRA consistently outperforms decoder tuning in leaf counting and disease classification, with their method achieving high performance in both tasks, approaching the results of SOTA models. Zanella et al. [48] explored few-shot adaptation of Vision Language Models (VLMs) using LoRA, showing that their CLIP-based method not only achieves substantial improvements but also reduces training times.
@@
The application of foundation models in biometrics is still very limited, with only few recent works started investigating their potential. For example, Iris-SAM [20], which is based on the foundation model SAM [28], fine-tunes it on ocular images for iris segmentation, achieving an average segmentation accuracy that surpasses the best baseline by a substantial margin of 10% on the ND-IRIS-0405 dataset. Arc2Face [34] is an identity-conditioned face foundation model that generates photo-realistic images based on the ArcFace embedding of a person. To showcase the performance of the generated data, they train a FR model on synthetic images from their model, achieving superior performance compared to existing synthetic datasets [7]. Recognizing the immense potential of foundation models across diverse tasks, this study breaks new ground by exploring their adaptation for face recognition, a path, to the best of our knowledge, unexplored until now.
@@ 3 Methodology
This section presents our approach FRoundation, for optimizing vision foundation models for face recognition. This section starts by providing details on the selected baseline foundation models, CLIP [35] and DINOv2 [33]. Then, we provide details on the adaptation mechanism used to optimize foundation models for downstream tasks. Finally, we conclude by describing the extension of pretrained foundation models with LoRA for face recognition.
@@ 3.1 Baseline Foundation Models
We selected two state-of-the-art foundation models, CLIP and DINOv2, to conduct our studies in this paper. These models proved to be generalizable across different downstream tasks [33, 35] and achieved very competitive results with zero-shot learning. Previous works [49, 15] also proved that fine-tuning these models using, for example LoRA, could lead to SOTA performance on specific downstream tasks.
@@ 3.2 Fine-tuning with LoRA
In this work, we utilized Low-rank adaptation (LoRA) [24] to fine-tune the considered foundation model. LoRA was initially developed to fine-tune LLMs for specific downstream tasks, but it can be applied to any neural network with dense layers. Its development was inspired by [1], which demonstrates that a low-dimensional reparameterization can be as effective for fine-tuning as the full parameter space. This indicates that pre-trained models actually reside on a low intrinsic dimension. Building on this concept, LoRA freezes the pre-trained model weights and inserts trainable rank decomposition matrices into each layer of pretrained transformer architecture. For a pre-trained weight matrix , LoRA utilizes a low-rank decomposition to restrict its update by where and with the rank . does not receive gradient updates during the training process, while only and are updated. When fine-tuning, this approach significantly reduces the number of trainable parameters compared to fine-tuning all model parameters, while also not introducing any additional inference latency. The latter is achieved by computing . After fine-tuning the model, the adapter weights are merged with the base model weight to compute the final model weights .
During the forward pass, the low-rank matrix product is scaled by where is a constant. When optimizing with Adam [27], tuning is roughly the same as tuning the learning rate [24]. This scaling factor causes gradient collapse as the rank increases, resulting in larger ranks performing no better than smaller ones [26]. To tackle this issue, the rank-stabilized LoRA (rsLoRA) [26] method proposes to scale the low-rank matrix with . Gradients do not collapse with rsLoRA, and training with higher ranks has been experimentally validated to improve performance. This method allows for an effective compute/performance trade-off, where higher ranks can be used to achieve higher performance at the cost of increased training compute.
@@ 3.3 FRoundation
Utilizing pretrained foundation model for face recognition leads to suboptimal results, as will be presented in Section 5.1. Thus, we propose to fine-tune the considered foundation models using LoRA. This requires extending the pretrained ViT models with LoRA layers before fine-tuning them on dedicated datasets. It is possible to apply LoRA to the , , , and projection layers, which refer respectively to the query, key, value, and output projection matrices in the self-attention module of ViT architecture [19]. Following [24], we adapt the weight matrices of q and v only, as it was shown that adapting their respective weights and yields the best results on different downstream tasks [24].
The transformer encoder consists of alternating layers of Multiheaded Self-Attention (MSA) and multilayer perceptron (MLP) blocks. Layer Normalization (LN) is applied before every block and residual connections after every block. LoRA is applied exclusively to the attention weights, leaving the MLP unchanged for both simplicity and parameter efficiency [24]. As illustrated in Fig 1, given a facial image, the image is divided into non-overlapping patches which are then mapped into patch embeddings using a linear projection layer. Additionally, a learnable embedding known as the class (CLS) token [19] is appended to the sequence of embedded patches. The goal of the CLS token is to serve as the image representation, which we utilized to obtain feature representation from input face samples. Then, position embeddings are incorporated into the patch embeddings to preserve positional information. The resulting sequence of embedding vectors is used as input for the encoder. In MSA, we run heads in parallel, each with its own set of , , and . In each head, LoRA layers operate separately and have their own distinct weights. As shown in Fig 2, for an embedding , the computation of the , , and projection layers in head are:
@@
The output of the projection layer is then passed to the MLP, completing the execution of a single transformer block. transformer layers are used to transform the image tokens into feature representations where denotes the output of the -th transformer block. Each -th transformer layer processes the output vectors of the previous layer. The final hidden state of the CLS token is employed as feature representation. To optimize the considered foundation models for FR, we proposed to fine-tune the models with the widely adapted margin-penalty softmax loss for face recognition [42, 17]. Specifically, we extended the architecture with an additional multi-class classification layer and utilized CosFace loss [42] as a margin-penalty softmax loss.
| Method | Backbone | LFW | CFPFP | AgeDB30 | CALFW | CPLFW | Avg. | IJB-B | IJB-C | ||||

### v3

@@ Abstract
Foundation models are predominantly trained in an unsupervised or self-supervised manner on highly diverse and large-scale datasets, making them broadly applicable to various downstream tasks. In this work, we investigate for the first time whether such models are suitable for the specific domain of face recognition (FR). We further propose and demonstrate the adaptation of these models for FR across different levels of data availability, including synthetic data. Extensive experiments are conducted on multiple foundation models and datasets of varying scales for training and fine-tuning, with evaluation on a wide range of benchmarks. Our results indicate that, despite their versatility, pre-trained foundation models tend to underperform in FR in comparison with similar architectures trained specifically for this task. However, fine-tuning foundation models yields promising results, often surpassing models trained from scratch, particularly when training data is limited. For example, after fine-tuning only on 1K identities, DINOv2 ViT-S achieved average verification accuracy on LFW, CALFW, CPLFW, CFP-FP, and AgeDB30 benchmarks of 87.10%, compared to 64.70% achieved by the same model and without fine-tuning. While training the same model architecture, ViT-S, from scratch on 1k identities reached 69.96%. With access to larger-scale FR training datasets, these performances reach 96.03% and 95.59% for the DINOv2 and CLIP ViT-L models, respectively. In comparison to the ViT-based architectures trained from scratch for FR, fine-tuned same architectures of foundation models achieve similar performance while requiring lower training computational costs and not relying on the assumption of extensive data availability. We further demonstrated the use of synthetic face data, showing improved performances over both pre-trained foundation and ViT models. Additionally, we examine demographic biases, noting slightly higher biases in certain settings when using foundation models compared to models trained from scratch. We release our code and pre-trained models’ weights at github.com/TaharChettaoui/FRoundation.
@@ 1 Introduction
Significant progress has been made in developing foundation models trained on extensive and diverse datasets recently [47, 3, 28]. Once these models are trained, they can be adapted to a wide array of downstream tasks, providing a versatile basis. Inspired by the success of large language models (LLM) [22, 10, 57, 16, 2], similar large-scale foundation models have been explored for various perception tasks [51, 47, 3, 28]. Often based on the ViT architecture [23], which has been shown to match or exceed the performance of traditional methods in large-scale image classification tasks, visual foundation models are becoming increasingly popular due to their strong generalization capabilities when trained on larger datasets. These models can be optimized to perform template extraction, using zero or few-shot learning approaches, making them highly versatile for biometric applications, where collecting a large set of training data, e.g. face images, is technically and legally challenging. Although foundation models hold considerable promise for a wide range of applications, their adaptation for face recognition (FR) has not been explored in previous research, which motivates this study. In this work, we utilize visual foundation models to enhance performance for the downstream task of FR while minimizing reliance on extensive amounts of data, leveraging their pre-training on diverse datasets.
@@
To explore the potential of foundation models in FR, we evaluate the performance of different versions of two widely used foundation models, namely DINOv2 [47] and CLIP [51]. Although foundation models demonstrate strong generalization capabilities, our experiments show that they perform poorly compared to state-of-the-art (SOTA) FR models [20, 60, 5, 19] on various benchmarks. To enhance their effectiveness for the downstream task of FR, we propose to fine-tune the considered foundation models using low-rank adaptation (LoRA) [30]. LoRA integrates trainable low-rank decomposition matrices into each transformer block while keeping the pre-trained model weights frozen. After fine-tuning on dedicated datasets, the LoRA layers adapt to the downstream task of FR, resulting in a significant increase in performance. For example, the smallest released pre-trained models of DINOv2 and CLIP achieve average verification accuracies of 64.70% and 82.64%, respectively. After fine-tuning on CASIA-WebFace [66], their accuracies increase to 90.94% and 92.13%, respectively.
In addition to examining the performance of fine-tuned foundation models for FR, we compare them to a Vision Transformer (ViT) trained from scratch on different subsets of a small-scale training dataset, namely CASIA-WebFace. The goal of this experiment is to leverage the versatility of foundation models, which are trained on extensive and diverse datasets, making them strong candidates for small-scale fine-tuning, as they benefit from the data on which they were previously trained. Additionally, we aim to address the challenges in data collection for the downstream task of FR, which can be tedious due to legal privacy regulations requiring strict consent, ethical considerations regarding individual rights, and technical limitations in collecting large and diverse training datasets. Our experimental results demonstrate that, under conditions of low data availability, fine-tuned foundation models significantly outperform models trained from scratch, leveraging their pre-trained knowledge. On the other hand, as data availability increases, the performance of models trained from scratch becomes competitive. To provide insight into foundation models performances when large-scale finetuning datasets are available and to provide comparison results to recent SOTA FR models, we finetuned/trained different instances of CLIP, DINOv2, and ViT on two large-scale datasets, MS1MV2 and WebFace4M, that are widely used in the literature. Our results validated that models trained from scratch can eventually compete with or outperform fine-tuned foundation models when using large-scale datasets, highlighting the critical importance of selecting the appropriate training strategy based on dataset size.
With the growing use of synthetic data to develop FR models [9, 44, 48], which enables privacy-preserving training, we explore the performance of these solutions and demonstrate that foundation models outperform models trained from scratch when both are trained on the same synthetic data. The presented models in this paper are evaluated under cross-validation settings on mainstream challenging benchmarks, including ones with cross-age, cross-pose, and large-scale verification benchmarks. These evaluations are also enriched with demographic bias evaluation on racial face in the wild (RFW) by reporting verification accuracies as well as bias assessing metrics, e.g. skewed error ratio (SER) and standard deviations (STD). Our results on RFW, which are aligned with previous works [63, 61, 44, 25, 62], report demographic bias in deep learning-based FR models.
@@ 2 Related Work
Recent advances in FR performance have been primarily driven by advancing development in neural network architectures [29, 23], innovations in training loss functions [60], and the availability of large-scale training datasets [27, 75, 11]. The majority of recent FR studies [60, 20, 45, 5, 32, 34] employ ResNet-like [29] architectures, with some recent works [56, 36] exploring ViT-based [23] architectures. Most of these studies [60, 20, 45, 5, 32, 34] focus on innovations in training loss functions. FR training losses can be broadly categorized into two groups: metric learning (e.g., Triplet loss [53]) and multi-class classification learning (e.g., Softmax loss [40]). Metric learning losses [55, 53], such as Triplet loss, explicitly encourage the model to learn discriminative feature representations by minimizing distances (e.g., Euclidean distance) between samples of the same label while maximizing distances between samples of different labels. In contrast, margin-penalty softmax loss employs cross-entropy over a softmax layer to implicitly guide the model in learning discriminative features. This is achieved by deploying a margin penalty on the angle or cosine angle between samples and their respective class centers, aiming at pushing samples closer to their respective class centers and further away from other class centers. Innovations in margin-penalty softmax loss have focused on the geometric deployment of penalty margins, whether fixed [60, 20] or adaptive [45, 5, 32, 34], yielding significant recognition improvements on mainstream benchmarks. These advancements in FR research have been made possible by large-scale training datasets [27, 75, 11], which enable the training of deep networks with millions of parameters. Notable datasets include CASIA-WebFace [66], MS-Celeb-1M [27] and its subsets (MS1MV2 [20] and MS1MV3 [21]), VGGFace2 [11], and WebFace260M [75] with its subsets (WebFace42M, WebFace12M, and WebFace4M). However, these datasets are collected from the internet, raising discussion about the ethical use of such data in FR development [9]. Such concerns, combined with the technical challenges of assembling large-scale datasets, have motivated researchers [65, 35, 7, 6] to explore the use of synthetically generated data in FR development. These challenges in collecting large-scale face datasets are among the key motivations for this work. As detailed later in this paper, we studied the impact of fine-tuning dataset size on the foundation model performances, providing insights into the model performance when extremely small subsets are available (e.g., 82k images of 1k identities) and in the case where large-scale datasets (e.g., 5.8M images of 85k identities) are accessible. We also opted, as detailed later in this paper, to fine-tune foundation models on synthetically generated data, exploring the potential of such data in adapting foundation models for the FR task.
Foundation models:
@@
Foundation models are defined by a substantial number of trainable parameters and are pre-trained on a large and diverse dataset, enabling them to adapt to a wide range of downstream tasks across various domains [4]. Vision foundation models are commonly structured around the use of Vision Transformers (ViTs) [23] and rely on self-supervised learning (SSL) [26]. SSL is a technique that trains models to learn representations from unlabeled data and is essential for training ViTs, which tend to perform poorly on small datasets [74]. Several high-performing vision foundation models specialized in various tasks have been released in recent years. Building on the success of the Segment Anything Model (SAM) [38], which introduced a foundation model for object segmentation, SAM2 [52] presents a model for real-time, promptable object segmentation in images and videos, achieving SOTA performance. CLIP [51] learns visual concepts from natural language supervision and can be applied to any visual classification by providing the names of the visual categories to be recognized. DINOv2 [47] foundation models generate universal features suitable for both image-level visual tasks, such as image classification and video understanding, as well as pixel-level visual tasks, including depth estimation and semantic segmentation.
@@
Vision foundation models are characterized by their generalization ability due to massive training data, but they tend to show poor performance when applied to domain-specific settings [59]. To adapt them to a downstream task, multiple approaches were considered in the literature. An example of that is the AdaptFormer [14] that replaces the MLP block in the transformer encoder with two identical MLP branches, where one mirrors the original network and the other introduces a lightweight module for task-specific fine-tuning, demonstrating significant improvements compared to fully fine-tuned models on downstream tasks. Another example is the ViT-Adapter [15] that proposes a method to adapt plain ViTs for dense prediction tasks by injecting image-related inductive biases into the ViT and reconstructing fine-grained multi-scale features, yielding SOTA results on COCO test-dev. Another approach is to insert trainable rank decomposition matrices, called Low-rank adaptation (LoRA) [30] layers while freezing the pre-trained model weights. In this work, we choose LoRA as our foundation model adapter, as recent studies [69, 18, 13, 68, 59] highlight the significant potential of LoRA for this purpose. For instance, integrating LoRA layers into DINOv2 has been successfully applied in the medical domain for two distinct tasks: capsule endoscopy diagnosis [69] and depth estimation in endoscopic surgery [18], both of which have proven superior performance in their respective fields. In another work on adapting foundation models for multiple plant phenotyping [13], Chen et al. demonstrated that LoRA consistently outperforms decoder tuning in leaf counting and disease classification, with their method achieving high performance in both tasks, approaching the results of SOTA models. Zanella et al. [68] explored few-shot adaptation of Vision Language Models (VLMs) using LoRA, showing that their CLIP-based method not only achieves substantial improvements but also reduces training times.
@@
The application of foundation models in biometrics is still very limited, with only a few recent works starting to investigate their potential. For example, Iris-SAM [24], which is based on the foundation model SAM [38], fine-tunes it on ocular images for iris segmentation, achieving an average segmentation accuracy that surpasses the best baseline by a substantial margin of 10% on the ND-IRIS-0405 dataset. Arc2Face [50] is an identity-conditioned face foundation model that generates photo-realistic images based on the ArcFace [20] embedding of a person. To showcase the performance of the generated data, they train a FR model on synthetic images from their model, achieving superior performance compared to existing synthetic datasets [9]. Recognizing the immense potential of foundation models across diverse tasks, this study uncovers new perspectives by exploring their adaptation for FR, a path that, to the best of our knowledge, has been unexplored until now.
@@ 3 Methodology
This section presents our approach FRoundation, for optimizing vision foundation models for FR. This section starts by providing details on the selected baseline foundation models, CLIP [51] and DINOv2 [47]. Then, we provide details on the adaptation mechanism used to optimize foundation models for downstream tasks. Finally, we conclude by describing the extension of pre-trained foundation models with LoRA for FR.
@@ 3.1 Baseline Foundation Models
We selected two SOTA foundation models, CLIP and DINOv2, to conduct our studies in this paper. These models proved to be generalizable across different downstream tasks [47, 51] and achieved very competitive results with zero-shot learning. Previous works [69, 18] also showed that fine-tuning these models using, for example, LoRA, could lead to SOTA performance on specific downstream tasks.
@@ 3.2 Fine-tuning with LoRA
In this work, we utilize Low-rank adaptation (LoRA) [30] to fine-tune the considered foundation model. LoRA was initially developed to fine-tune LLMs for specific downstream tasks, but it can be applied to any neural network with dense layers. Its development was inspired by [1], which demonstrates that a low-dimensional reparameterization can be as effective for fine-tuning as the full parameter space. This indicates that pre-trained models actually reside on a low intrinsic dimension. Building on this concept, LoRA freezes the pre-trained model weights and inserts trainable rank decomposition matrices into each layer of the pre-trained transformer architecture. For a pre-trained weight matrix , LoRA utilizes a low-rank decomposition to restrict its update by where and with the rank . does not receive gradient updates during the training process, while only and are updated. When fine-tuning, this approach significantly reduces the number of trainable parameters compared to fine-tuning all the model parameters, while also not introducing any additional inference latency. The latter is achieved by computing . After fine-tuning the model, the adapter weights are merged with the base model weight to compute the final model weights .
During the forward pass, the low-rank matrix product is scaled by where is a constant. When optimizing with Adam [37], tuning is roughly the same as tuning the learning rate [30]. This scaling factor causes gradient collapse as the rank increases, resulting in larger ranks performing no better than smaller ones [33]. To tackle this issue, the rank-stabilized LoRA (rsLoRA) [33] method proposes to scale the low-rank matrix with . Gradients do not collapse with rsLoRA, and training with higher ranks has been experimentally validated to improve performance. This method allows for an effective compute/performance trade-off, where higher ranks can be used to achieve higher performance at the cost of increased training computation.
@@ 3.3 FRoundation
Utilizing pre-trained foundation models for FR leads to suboptimal results, as will be presented in Section 5.1. Thus, we propose to fine-tune the considered foundation models using LoRA. This requires extending the pre-trained ViT models with LoRA layers before fine-tuning them on dedicated datasets. It is possible to apply LoRA to the , , , and projection layers, which refer respectively to the query, key, value, and output projection matrices in the self-attention module of ViT architecture [23]. Following [30], we adapt the weight matrices of and only, as it was shown that adapting their respective weights and yields the best results on different downstream tasks [30].
The transformer encoder consists of alternating layers of Multiheaded Self-Attention (MSA) and multilayer perceptron (MLP) blocks. Layer Normalization (LN) is applied before every block and residual connections after every block. LoRA is applied exclusively to the attention weights, leaving the MLP unchanged for both simplicity and parameter efficiency [30]. As illustrated in Fig 1, given a facial image, the image is divided into non-overlapping patches which are then mapped into patch embeddings using a linear projection layer. Additionally, a learnable embedding known as the class (CLS) token [23] is appended to the sequence of embedded patches. The goal of the CLS token is to serve as the image representation, which we utilize to obtain feature representation from input face samples. Then, position embeddings are incorporated into the patch embeddings to preserve positional information. The resulting sequence of embedding vectors is used as input for the encoder. In MSA, we run heads in parallel, each with its own set of , , and . In each head, LoRA layers operate separately and have their own distinct weights. As shown in Fig 2, for an embedding , the computation of the , , and projection layers in head are:
@@
The output of the projection layer is then passed to the MLP, completing the execution of a single transformer block. transformer layers are used to transform the image tokens into feature representations where denotes the output of the -th transformer block. Each -th transformer layer processes the output vectors of the previous layer. The final hidden state of the CLS token is employed as feature representation. To optimize the considered foundation models for FR, we proposed to fine-tune the models with the widely adapted margin-penalty softmax loss for FR [60, 20]. Specifically, we extended the architecture with an additional multi-class classification layer and utilized CosFace loss [60] as a margin-penalty softmax loss.
| Method | Backbone | LFW | CFP-FP | AgeDB30 | CALFW | CPLFW | Avg. | IJB-B | IJB-C | ||||